# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v5

@@ 6.2 Main Results
Figure 4(a) shows no significant correlation between the public data ratio and privacy-preserving capability in the same training round. In addition, only using local data has a higher risk of privacy leakage than augmented methods. Additionally, Figure 4(b) shows $\rho$ trends per round. Initially, the $\rho$ for augmented settings increases, then decreases or converges, while the base-data-only scores continue to rise, especially in the code domain. This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
Scalability.
We scale the client number to 100 and 2 clients are randomly selected for FedDIT in each round. As the number of clients increases, the amount of local data on each client gradually grows. Table 4 shows that the model trained on only base data even outperforms Random Sampling in domains other than code and narrows the gap with FedDCA. However, because FedDCA aims to maximize the cross-client domain coverage, it achieves higher domain coverage and better performance.
Data Selection Scenarios.
We conduct the experiments on two settings: A) For each domain’s FedDIT, we randomly dispatch the in-domain data equally to each client. For example, in the code domain, as the size of the public dataset is 20,022, each client is assigned about 2,000 samples. Then we perform FedDIT on each client’s whole data. B) Following A’s setting, we first perform FedDCA to determine the client center set, and then we conduct the data selection based on the client center set through dense retrieval to select 200 samples for each client. Finally, we perform FedDIT on the selected data. Figure 6(b) demonstrates that FedDCA can achieve comparable performance to training with the full dataset by using only 10% data through data selection.
Held-out Setting.
Considering the held-out setting in the financial domain, given that the training set FinGPT and the test sets FPB, FiQA, and TFNS are all related to sentiment analysis tasks. We keep the setting of test sets and replace the FinGPT’s instructions in public data with data from a financial QA dataset (Sujet). The clients’ local data are still randomly sampled from FinGPT. This approach yields held-out public data.
Figures 6(d), 6(e) and 6(f) shows, FedDCA still achieves performance improvements compared to other baselines for the held-out setting. Additionally, using the Random Sampling data augmentation strategy resulted in performance degradation on the FiQA dataset. This further underscores the necessity of selecting an appropriate data augmentation strategy.
6.3 Ablation Study
Effect of Retrieval Number.
We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 5(b). We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance except for the code domain. Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
Similarity Threshold.
Figure 5(a) shows the performance and the domain coverage of FedDCA w and w/o similarity threshold $\alpha$, where FedDCA w/o $\alpha$ is marked as FedDCA. Results show that FedDCA is slightly better than FedDCA in code, financial, and mathematical domains.
Impact of Different Cluster Number.
The hyperparameter $\xi$ is the number of clusters in the k-means algorithm. The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients. We report the domain coverage of the augmented dataset via FedDCA with different $\xi$ on the four domains in Figure 6(c). Results show that there is no best $\xi$ for all domains and $\xi=N$ is usually an acceptable choice.
+ FedDCA H-Eval M-Med FPB GSM8K
Random Sampling 32.51 (-0.42) 71.80 (+0.50) 63.92 (-0.27) 47.82 (+0.44)
Direct Retrieval 33.89 (-0.25) 72.10 (-0.10) 67.78 (+1.47) 52.53 (+1.66)
LESS 28.32 (+0.28) 70.80 (-0.20) 61.03 (+0.47) 43.92 (+0.79)
Self-Instruct 33.15 (+0.23) 70.70 (-1.20) 60.21 (+0.48) 50.25 (-0.54)
Table 3: Plug-and-Play. Performance (%) of FedDCA with different data augmentation methods. FedDCA can match full dataset performance using just 10% of the data through selection.
Metric Zero-shot Base Data Random Sampling FedDCA
H-Eval 29.88/- 34.14/0.8282 34.75/0.8685 35.92/0.9242
M-Med 70.60/- 72.40/0.8377 69.90/0.8497 73.30/0.9090
FPB 55.94/- 66.74/0.9339 61.05/0.9408 67.16/0.9800
GSM8K 23.27/- 49.12/0.8709 47.23/0.8812 50.26/0.9118
(a) W and w/o $\alpha$
(b) Different Retrieval Amounts
Figure 5: Solid bars and lines are performance. Performance (%) and domain coverage of FedDCA (a) w and w/o similarity threshold $\alpha$; (b) w different retrieval amounts. Experiments show the necessity of similarity threshold for amplifying diversity and the strong applicability of FedDCA with various augmentation scale.
(a) Computational Cost
(b) Data Selection Scenarios
(c) Different $\xi$
(d) Held-out. FPB
(e) Held-out. FiQA
(f) Held-out. TFNS
Figure 6: (a) Computational cost comparison. (b) Domain coverage for different $\xi$. (c) Data selection scenarios. (d-f) Performance in the held-out setting. Results show the efficiency of FedDCA, the effectiveness of FedDCA in data selection scenarios, and the robustness of FedDCA in the held-out setting.
7 Conclusion
We reveal that in FedDIT, data heterogeneity has a non-monotonic relationship with model performance, while cross-client domain coverage significantly affects model effectiveness. To address this, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based instruction augmentation. The selection process ensures comprehensive cross-client domain coverage by greedily picking centers that represent maximal distinct regions of the instruction space. FedDCA further reduces client-side computation and enhances scalability using heterogeneous encoders. Extensive experiments comparing FedDCA with other nine baselines and across four domains demonstrate its effectiveness and plug-and-play capability.
Appendix
Appendix A Methodology
A.1 Complementary Description
Notation Table.
For convenience, we provide a detailed notation table in Table 5.
NOTATIONS REMARK
$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.
Table 5: Frequently used notation.
A.2 Direct Domain Data Retrieval
In this section, we first show the detail of instruction-based dense retrieval, which is both used in direct retrieval and FedDCA. Then, we explain the algorithm of the method direct retrieval.
A.2.1 Instruction Based Dense Retrieval
[…]
A.2.2 Direct Retrieval
The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation. The detailed algorithm is described in Algorithm 2. For each client, we start by encoding the local instructions using the encoder model $w_{enc}$. Next, we apply the k-means algorithm to cluster the embeddings into $\xi$ clusters. Then $\xi$ cluster centers are sent to the server for retrieval-based domain data augmentation. Subsequently, the retrieved public data is sent to the client for instruction tuning.
Algorithm 2 Direct domain data retrieval
1:Parameters:
2:Clients’ local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Local datasets $D^{l}=\{D^{l}_{1},D^{l}_{2},\ldots,D^{l}_{N}\}$; Number of clusters $\xi$; Encoder model $w_{enc}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
3:$D^{g}\leftarrow\emptyset$
4:for $i\in\{1,2,\ldots,N\}$ do
5:     $\mathcal{E^{\prime}}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$triangleright Encode the local instructions
6:     $\mathcal{C}\leftarrow\text{k-means}(\xi).fit(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}$
7:     $\mathcal{S}\leftarrow\mathcal{C}\cdot\mathcal{E}^{T}$triangleright Compute the similarity score between $\mathcal{C}$ and $\mathcal{E}$
8:     $\mathcal{I}\leftarrow\emptyset$
9:     for $j\in\{1,2,\ldots,\xi\}$ do
10:         $\mathcal{S}^{\prime}\leftarrow\{s\mid s\in S_{j},\text{index}(s)\notin\mathcal{I}\}$triangleright Filter the selected indices
11:         $\mathcal{I^{\prime}}\leftarrow\text{indices of the top-}\frac{N_{k}^{p}}{\xi}\text{ elements in }\mathcal{S}^{\prime}$triangleright Retrieve the top $\frac{N_{k}^{p}}{\xi}$ indices
12:         $\mathcal{I}\leftarrow\mathcal{I}\cup\mathcal{I^{\prime}}$triangleright Update the selected indices
13:     end for
14:     $D^{g}_{i}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$triangleright Selected public instructions
15:     $D_{i}\leftarrow D^{g}_{i}\cup D_{i}^{l}$triangleright Obtain the augmented instruction dataset $D_{i}$
16:end for
Appendix B Related Works
Federated Instruction Tuning.
Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al. [2023c]; Wei et al. [2022]. Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks Ye et al. [2024b, a]; Zhang et al. [2023b] have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
PrivateLoRA Wang et al. [2023b] addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT Zhang et al. [2024b] focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
Non-IID data distribution is a common challenge in FL Karimireddy et al. [2020]; Li et al. [2020]; Sattler et al. [2019]; Li et al. [2019, 2021]. However, the impact of data heterogeneity on FedIT has not been fully explored. Shepherd Zhang et al. [2023b] constructing heterogeneity based on the topic of general instruction datasets. Results demonstrate that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT. To fully understand the impact of data heterogeneity on FedDIT, we conduct a comprehensive analysis in this work and unveil a crucial finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT (Section 3).
Domain Instruction Augmentation.
In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al. [2023]; Luo et al. [2024], medical Zhang et al. [2023d], financial Yang et al. [2023b, a]; Zhang et al. [2023a]; Wu et al. [2023], mathematical Yue et al. [2024]; Luo et al. [2023]). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets Wang et al. [2023a]; Zhang et al. [2023e]; Jiao et al. [2023]; Lee et al. [2024]; Xia et al. [2024]. For instance, Parrot Jiao et al. [2023] enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA Lee et al. [2024] and LESS Xia et al. [2024] attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA Lee et al. [2024] uses instructions without responses for effective retrieval. LESS Xia et al. [2024] assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct Luo et al. [2024]; Wan et al. [2023]: Explore-Instruct Wan et al. [2023], for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works Yue et al. [2024]; Zhou et al. [2024] highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2 Yue et al. [2024] firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0 Zhou et al. [2024] distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
Appendix C Experiments
C.1 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means Wu [2012], we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning Khosla et al. [2020]. We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
Projector.
To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA. We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side. As shown in Figure 7, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
Figure 7: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
Prompts Used in the Self-Instruct Data Generation.
To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 8. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
Prompt Used for Memory Extraction Attack.
As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 9, which is exactly the prefix of the Alpaca’s template.
Figure 9: Prompt used for memory extraction attack.
C.2 Baselines
Method Privacy Preserving API Cost Additional Information Performance Degradation Domain Coverage Oriented
FedAvg McMahan et al. [2017]
Random Sampling
LESS Xia et al. [2024]
Self-Instruct Wang et al. [2022]
Direct Retrieval
FedDCA (ours)
In Table 6, we highlight five critical aspects of our approach compared with other methods Xia et al. [2024]; Wang et al. [2022]; Zhang et al. [2024b]. Our comparison focuses on five key dimensions:
  • Privacy Preserving: Most methods except Self-Instruct maintain data privacy, we analyze the privacy-preserving ability of FedDCA in Sections 5 and C.3.
  • API Cost: FedDCA requires no api calls. As Self-Instruct utilizes the GPT to generate instructions, it requires additional API calls, which incurs costs.
  • Additional Information: LESS needs validation set access for gradient-based retrieval, whereas FedDCA works with only the basic client’s local training data.
  • Performance Degradation: From Table 2 we see that there occurs performance degradation for FedAvg, Random Sampling, and Self-Instruct, while FedDCA avoids this issue.
  • Domain Coverage: Uniquely, FedDCA is specifically designed to optimize domain coverage for distributed environments, a feature absent in other baseline methods.
Specifically, Self-Instruct Wang et al. [2022] represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT Zhang et al. [2024b] attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training. However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly regarding both time and computational resources. Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario. LESS Xia et al. [2024] represents the augmentation methods through gradient feature retrieval. However, its performance is underwhelming. Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data. This process becomes computationally burdensome with large public datasets.
In conclusion, these comparisons demonstrate FedDCA’s distinctive approach to balancing privacy preservation, computational overhead and domain coverage optimization, while maintaining minimal additional requirements.
C.3 Domain Inference Attacks
@@
Example 1: Both $w_{1}$ and $w_{2}$ take "hello world" as input, and the cosine similarity between their embeddings is 0.1829.
@@
C.4 How Far can FedDCA Go
Assume an extreme setting, where exists a pre-trained model that has already fine-tuned on the whole in-domain data of the internet or even all open-source public instruction datasets (it is actually unrealistic or temporary, as new data is constantly being generated), is there any room for instruction augmentation or FedDCA? To answer this question, we conduct the following experiments on four settings: A) Fine-tuning on the whole public dataset. B) Following A, perform further federated fine-tuning with only private clients’ local dataset. C) Following A, each client randomly samples 100 public data for fine-tuning. D) Following A, each client samples 100 public data through FedDCA and then performs FedDIT.
Setting H-Eval M-Med FPB FiQA TFNS GSM8K
A 38.41 69.10 75.33 41.09 71.60 53.75
B 43.90 64.30 77.97 66.54 78.26 56.40
C 41.46 66.90 74.50 29.81 70.51 57.01
D 44.12 70.90 81.43 72.00 78.81 58.90
Table 7: Performance (%) of setting A, B, C, and D in each domain.
As shown in Table 7, even with only 100 samples selected via FedDCA, the method not only prevents performance degradation but also helps the model achieve better generalization within the domain. This is attributed to exposure to more diverse in-domain data during training. Meanwhile, it can be observed that random sampling still leads to performance degradation compared to Setting A, and even performs worse than Setting B, where no data augmentation is applied. In the financial domain, performance degradation is also evident.
In conclusion, this experiment further highlights the necessity and effectiveness of FedDCA.
C.5 Augmentation Strategy Visualization

### v6

@@ B.5 Privacy Analysis
Figure˜7(a) shows no significant correlation between the public data ratio and privacy-preserving capability in the same training round. In addition, only using local data has a higher risk of privacy leakage than augmented methods. Additionally, Figure˜7(b) shows the trends of the average ROUGE-L score per round. Initially, the average ROUGE-L score for augmented settings increases, then decreases or converges, while the base-data-only scores continue to rise, especially in the code domain. This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
Domain Inference Attack.
@@
Example 1: Both $w_{1}$ and $w_{2}$ take ‘‘hello world’’ as input, and the cosine similarity between their embeddings is 0.1829.
@@
MMLU-Med FPB GSM8K
Zero-shot 70.60/- 55.94/- 23.27/-
Base Data 72.40/0.8377 66.74/0.9339 49.12/0.8709
Random Sampling 69.90/0.8497 61.05/0.9408 47.23/0.8812
FedDCA 73.30/0.9090 67.16/0.9800 50.26/0.9118
B.6 Scalability
(a) FPB
(b) FiQA
(c) TFNS
Figure 8: Held-out setting. We report the performance of FedDCA and other baselines for (a) FPB, (b) FiQA, and (c) TFNS datasets. Results show the robustness of FedDCA in the held-out setting.
Figure 9: Data selection scenarios. We compare the performance of using the full dataset (Setting A) and using only 10% of data selected through FedDCA (Setting B), demonstrating FedDCA’s efficiency in data utilization.
Experimental Setup.
To evaluate the scalability of FedDCA, we conduct experiments with 100 clients, where 2 clients are randomly selected for FedDIT in each round. As the number of clients increases, the amount of local data on each client gradually grows, allowing us to assess how FedDCA performs in larger-scale distributed settings.
Results.
Table˜6 presents the performance and domain coverage across different domains. Interestingly, we observe that the model trained on only base data even outperforms Random Sampling in domains other than code and narrows the gap with FedDCA. This suggests that with sufficient local data, the benefits of data augmentation become more nuanced. However, FedDCA maintains its advantage by consistently achieving higher domain coverage and better performance across all domains, demonstrating its effectiveness in large-scale federated settings. This superior performance can be attributed to FedDCA’s strategic approach to maximizing cross-client domain coverage, which remains crucial even when dealing with a larger overall volume of local data distributed across a greater number of clients.
B.7 Data Selection Scenarios
Experimental Setup.
We conduct the experiments on two settings: A) For each domain’s FedDIT, we randomly dispatch the in-domain data equally to each client. For example, in the code domain, as the size of the public dataset is 20,022, each client is assigned about 2,000 samples. Then we perform FedDIT on each client’s whole data. B) Following A’s setting, we first perform FedDCA to determine the client center set, and then we conduct the data selection based on the client center set through dense retrieval to select 200 samples for each client. Finally, we perform FedDIT on the selected data.
Results.
Figure˜9 demonstrates that FedDCA can achieve comparable performance to training with the full dataset by using only 10% data through data selection. This efficiency in data utilization highlights the effectiveness of FedDCA’s data selection strategy, which maintains model performance while significantly reducing the required training data.
B.8 Held-out Setting
Considering the held-out setting in the financial domain, given that the training set FinGPT and the test sets FPB, FiQA, and TFNS are all related to sentiment analysis tasks. We keep the setting of test sets and replace the FinGPT’s instructions in public data with data from a financial QA dataset (Sujet-Finance-Instruct-177k [^4] https://huggingface.co/datasets/sujet/Sujet-Finance-Instruct-177k). The clients’ local data are still randomly sampled from FinGPT. This approach yields held-out public data.
Figures˜8(a), 8(b) and 8(c) shows, FedDCA still achieves performance improvements compared to other baselines for the held-out setting. Additionally, using the Random Sampling data augmentation strategy resulted in performance degradation on the FiQA dataset. This further underscores the necessity of selecting an appropriate data augmentation strategy.
B.9 Empirical Tightness of FedDCA’s Greedy Approximation
FedDCA’s optimization objective is equivalent to the classic sub-modular facility-location problem, which guarantees that the simple greedy algorithm attains at least a 11/e1-1/e fraction of the optimal value in the worst case. To gauge how tight this bound is in practice, we benchmark the FedDCA algorithm against increasingly stronger (but more expensive) baselines on three domains: Medical, Financial and Mathematical. In particular we position FedDCA between the beam search (with various beam widths) and an infeasible brute-force solver that attempts to enumerate all candidate subsets. The beam search, which explores multiple promising branches simultaneously, serves as a realistic upper bound that is strictly better than greedy yet still tractable.
Experimental Setup.
  1. 1.
    Metric: Domain coverage $d(D^{d},\mathcal{P})$ where $\mathcal{P}$ is the selected $N$ client centers, defined as Eq. 2.
  2. 2.
    Baselines: $\mathcal{C}_{\text{all}}$ (default FedDCA) selects client centers using the union of each client’s uploaded cluster center set $\mathcal{C}_{\text{all}}$ as the selection reference to compute marginal gains; $\mathcal{P}^{*}_{C}$ is the unreachable optimum under the same selection reference $\mathcal{C}_{\text{all}}$—marked timeout, requiring evaluation of $\binom{100}{10}\approx 1.73\times 10^{13}$ distinct candidate sets; $D^{d}$ uses $D^{d}$ as the selection reference to compute marginal gains, which is often impractical in real-world federated settings (please refer to Section˜5.2); $\mathcal{P}^{*}_{D}$ is the optimum under $D^{d}$—also timeout; Beam search with widths $256,512,1024,2048$ and, for Medical, $9182$ (does not exceed the baseline $D^{d}$ until width $9182$).
  3. 3.
    Approximation ratio: We report
    $\frac{d(D^{d},\mathcal{P}_{\mathcal{C}_{\text{all}}})}{\max_{w}d(D^{d},\mathcal{P}_{\text{beam-}w})}\times 100\%,$ (10)
    where the denominator is the best coverage attained by the beam search. $\mathcal{P}_{\mathcal{C}_{\text{all}}}$ is the selected client center set by FedDCA using $\mathcal{C}_{\text{all}}$ as the selection reference.
Domain $\mathcal{C}_{\text{all}}$ $\mathcal{P}^{*}_{C}$ $D^{d}$ $\mathcal{P}^{*}_{D}$ Beam-256 Beam-512 Beam-1024 Beam-2048 Beam-9182 Approx. Ratio
Med. 0.6772 timeout 0.6782 timeout 0.6776 0.6776 0.6776 0.6778 0.6782 99.85%
Fin. 0.8379 timeout 0.8418 timeout 0.8422 0.8427 0.8427 0.8427 99.43%
Math. 0.6391 timeout 0.6699 timeout 0.7401 0.7401 0.7404 0.7405 86.30%
Table 7: Empirical tightness analysis of FedDCA’s greedy approximation. Both theoretical optima ($\mathcal{P}^{*}_{C}$ and $\mathcal{P}^{*}_{D}$) are computationally unattainable (timeout). The approximation ratio is computed against the best beam search result.
Results.
As shown in Table˜7, FedDCA using $\mathcal{C}_{\text{all}}$ already achieves nearly optimal coverage in Medical and Financial domains, within 99.85% and 99.43% of an extensive beam search respectively, and still attains a respectable 86.30% in the more challenging Mathematical domain.
Note that both theoretical optima ($\mathcal{C}_{\text{all}}$ and $D^{d}$ as reference) are computationally unattainable (timeout), so the beam search constitutes the strongest practical competitor.
In conclusion, these real-world ratios vastly exceed the theoretical lower bound of $1-1/e\approx 63\%$, further corroborating that the simple greedy procedure is both efficient and effectively near-optimal for real-world federated deployments.
Appendix C Robustness to Data Heterogeneity
The efficacy of retrieval-based augmentation strategies in FedDIT, such as Direct Retrieval and our proposed FedDCA, is inherently linked to the semantic characteristics of clients’ local private data, as these inform the selection of public instructions. A critical consideration is how varying degrees of inter-client data heterogeneity impact the ability to construct diverse and comprehensively covering augmented datasets. Intuitively, if clients possess highly similar local data, methods relying on these local signals might struggle to achieve broad cross-client domain coverage and instruction diversity in the augmented sets, which are crucial for a well-generalized global model.
This section, therefore, investigates the robustness of FedDCA, Direct Retrieval, and Random Sampling to a spectrum of inter-client data heterogeneity. We focus on these methods as their augmentation processes are either directly influenced by local data distributions (FedDCA, Direct Retrieval) or serve as a data-agnostic baseline (Random Sampling); other generative or complex retrieval methods often have different operational dependencies. We control heterogeneity using a Dirichlet distribution ($\text{Dir}(\beta)$) over semantic clusters of clients’ private instructions, where smaller $\beta$ values yield higher inter-client heterogeneity (more distinct local data) and larger $\beta$ values lead to lower heterogeneity (more similar local data). We evaluate performance across $\beta\in\{0.01,0.1,1.0,10\}$.
C.1 Experimental Setup
Inter-Client Data Heterogeneity Control.
We focus on the Financial domain, using FinGPT as the source for private instructions. First, we perform k-means clustering ($\xi=100$ clusters, consistent with Section˜4.1) on FinGPT instruction embeddings to create semantic pseudo-labels. To generate local data for 10 clients, each holding 100 instructions, we employ a Dirichlet distribution $\text{Dir}(\beta)$ over these clusters to define each client’s data composition, simulating varying degrees of inter-client data heterogeneity for $\beta\in\{0.01,0.1,1.0,10\}$.
Augmentation and Training.
Following the general setup in Section˜6.1, Random Sampling, Direct Retrieval (introduced in Section˜3), and FedDCA are used to obtain 1k augmented instructions per client before commencing FedDIT.
Evaluation Metrics.
Model performance is evaluated using FPB, FiQA, and TFNS accuracy. Augmented data characteristics are assessed using three key metrics:
1) Domain Coverage ($\uparrow$): As defined in Eq. 2, this measures how well the aggregated client data (local + augmented) $D^{c}$ covers the target financial domain $D^{d}$.
2) ICACS (Inter-Client Augmented Centroids Similarity $\downarrow$): For each client $k$, its augmented instruction embeddings are clustered (k-means, $\xi=10$) into centers $\mathcal{C}_{k}$. ICACS is the average pairwise cosine similarity between all such cluster centers across different clients. Lower values indicate more distinct augmented sets at a granular level.
3) RUAI (Ratio of Unique Augmented Instructions $\uparrow$): The ratio of unique instructions within the total cross-client augmented data pool $D^{c}$. Higher values signify less redundancy.
C.2 Results and Analysis
Dirichlet $\beta$ Augmentation Model Performance (%) Augmented Data Characteristics
Strategy FPB FiQA TFNS Domain Coverage ($\uparrow$) ICACS ($\downarrow$) RUAI ($\uparrow$)
0.01 Random Sampling 64.20 13.10 65.55 0.920 0.472 0.867
Direct Retrieval 67.20 21.50 69.00 0.938 0.434 0.909
FedDCA 67.90 36.30 74.40 0.985 0.405 0.943
0.1 Random Sampling 64.19 13.09 65.53 0.920 0.472 0.867
Direct Retrieval 66.31 19.11 67.62 0.929 0.441 0.882
FedDCA 67.24 35.27 73.32 0.982 0.411 0.938
1.0 Random Sampling 64.19 13.09 65.53 0.919 0.472 0.867
Direct Retrieval 65.50 17.80 66.79 0.924 0.448 0.879
FedDCA 66.94 34.82 72.66 0.977 0.414 0.927
10 Random Sampling 64.20 13.05 65.45 0.918 0.472 0.867
Direct Retrieval 64.37 15.20 66.10 0.922 0.452 0.871
FedDCA 66.73 34.51 72.23 0.972 0.423 0.921
Table 8: Performance (%), domain coverage, ICACS and RUAI across different levels of client data heterogeneity (controlled by Dirichlet parameter $\beta$) in the Financial Domain. Best model performance, domain coverage, ICACS and RUAI per $\beta$ setting are in bold.
Table˜8 systematically quantifies the impact of inter-client data heterogeneity on augmentation strategies. For small $\beta$ (i.e., high heterogeneity), the cross-client data exhibits substantial diversity, enabling Direct Retrieval to access a broader range of semantic regions in the public pool. This diversity translates into clear gains over Random Sampling in both performance (e.g., +6% FiQA at $\beta=0.01$) and coverage metrics. However, as $\beta$ increases and client data distributions become more homogeneous, the advantage of Direct Retrieval diminishes, with its performance and diversity metrics converging toward those of Random Sampling. This trend underscores that the effectiveness of Direct Retrieval is fundamentally driven by the diversity present in cross-client data.
Conversely, as inter-client heterogeneity decreases (i.e., local data distribution becomes more similar), Direct Retrieval’s effectiveness continuously degrades. Its model performance and domain coverage drop, showing a more sensitive effectiveness to the heterogeneity. Results indicate that with more homogeneous local data distributions, Direct Retrieval tends to fetch more similar public instructions due to the similarity between each client’s cluster centers.
Random Sampling exhibits stable, albeit consistently sub-optimal, performance and diversity metrics across all heterogeneity levels. This stability stems from its sampling strategy (oblivious to local data distributions) of the public dataset, which results in similar augmented data.
Additionally, FedDCA showcases strong robustness. It consistently achieves the best model performance across all tested $\beta$ values. While its diversity metrics (domain coverage, ICACS, RUAI) show minor sensitivity to the cross-client data distribution (optimally diverse at highest heterogeneity). This resilience stems from its coverage-oriented client center selection (Algorithm˜1), which effectively prevents redundant augmentation by selecting public instructions that add novel semantic aspects to the collective, even with similar cross-client data distributions. In conclusion, these findings underscore FedDCA’s robustness and adaptability irrespective of the specific degree of inter-client heterogeneity, makes it a more reliable and effective solution for real-world FedDIT deployments.
Appendix D Augmentation Strategy Visualization